OpenStudy (anonymous):

4(-x+4)=12 can someone help with that please

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

with what?

OpenStudy (anonymous):

math homework baby

OpenStudy (anonymous):

to the person who asked the question not you

OpenStudy (anonymous):

Okay its dealing with distribution. 4(-x + 4) distrubute that. 4x -x = -4x + (4x4) = 16 -4x + 16 =12 take 12 minus 16 = -4 -4x divided by -4 = X

OpenStudy (anonymous):

thank you very much guys

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

with what mfadym?

OpenStudy (anonymous):

with what?